The ratio of the length to the width of a rectangular advertising display is approximately 3.3 to 2.If the width of the display is 8 meters, what is the approximate length of the display in meters ?
Answer by jessica43(140)

To solve this you are going to want to set up equivalent fractions or ratios:
You know the ratio of the length to the width is 3.3 to 2, or 3.3/2.
You also know the length of the rectangle is x meters and the width of the rectangle is 8 meters, or x/8.
Now set these fractions equal to each other because they have ratios and solve for x:
(3.3/2)=(x/8)
1.65 = (x/8) (multiply each side by 8)
x=13.2
So the approximate length of the display is 13.2 meters.